# Matchings and loose cycles in the semirandom hypergraph model

### v1

@@ 5. Loose Hamilton cycles
The next steps are structurally the same as for matchings in Section 4. We will construct a directed bipartite graph on $A_{0}$ and $B_{0}$, but it will not have a perfect matching because a small number of nodes fail to have sufficiently high degree (and cause other nodes to “fail” as well). We deal with the failed nodes first, and in doing so add certain hyperedges, irrevocably, to our cycle-in-the-making $\mathcal{C}$. ($\mathcal{C}$ plays the role that $\mathcal{M}$ did for hypergraph matchings.) When the failed nodes are dealt with, most of our bipartite graph will remain, and what remains will have a perfect matching. The matching edges correspond to hyperedges in $\mathcal{H}$, and they complete $\mathcal{C}$ to a loose Hamilton cycle.
@@ 5.2. Phase 2
We introduce an additional structure, $\mathcal{P}$, central to the analysis. $\mathcal{P}$ is a graph on the odd points, consisting of vertex-disjoint paths. A path in $\mathcal{P}$ can be an isolated vertex. Initially, the edges in $\mathcal{P}$ are precisely the base pairs $B_{1}$. Eventually we will make $\mathcal{P}$ a single path and then a Hamilton cycle. Always, each edge in $\mathcal{P}$ is one of two types: it consists of the two odd points in a hyperedge $e=(b_{1},a,b_{2})\in\mathcal{H}$ (and all the apexes used for such edges will be distinct); or, as in the initial $\mathcal{P}$, it consists of a base pair in $B$ (and later it will be possible to perfectly match all these base pairs and the remaining apexes, in $D$). When $\mathcal{P}$ is a Hamilton cycle, this ensures that there is a corresponding loose hypergraph cycle $\mathcal{C}$ as desired.
In this phase we will again use a set $Q$ of nodes of $D$ “blocked” from use. Initially we will set $Q$ to consist of any nodes whose inedges include any double edge. Initially, $\mathcal{P}$ has $b_{x}$ components. Let us describe the phase just enough to count the number of nodes it will delete from each part of $D$, both to bound $\left|{Q}\right|$ and to confirm that, going in to Phase 3, $\left|{A}\right|=\left|{B}\right|$.
In Phase 2a, each action (see Figure 1) will assign an apex in $A_{X}$ to a hyperedge (along with two base points from different base pairs) and place it in $\mathcal{C}$, increasing the number of paths in $\mathcal{P}$ by 1, and deleting 2 base pairs from $B$. Phase 2a takes $a_{x}$ actions to resolve all the failed apexes, resulting in a total of $2a_{x}$ deletions from the set of base pairs in $B$, so that $\left|{B_{0}\setminus B}\right|=b_{x}+2a_{x}$, and $b_{x}+a_{x}$ components in $\mathcal{P}$. The part $A$ is unchanged, with $\left|{A_{0}\setminus A}\right|=a_{x}$.
@@
Let $S$ be the set of deleted nodes. For any given $\epsilon_{Q}>0$, let Lemma 9’s $\eta$ equal $\epsilon_{Q}/k$, and take the lemma’s corresponding $\delta$ to determine $2\epsilon_{D}$. Then, Lemma 9 ensures that the deleted set’s inneighbourhood has size $\left|{N^{-}(S)}\right|\leq(\epsilon_{Q}/k)n$. Since $D$ is $k$-out, the inneighbourhood’s outneighbourhood — which is to say, $Q$ — has size
$\left|{Q}\right|=\left|{N^{+}(N^{-}(S))}\right|\leq k(\epsilon_{Q}/k)n=\epsilon_{Q}n.$
@@ 5.2.1. Phase 2a
Upon offer of odd points $b_{1}$ and $b_{2}$, we propose a hyperedge $e=(b_{1},a,b_{2})$, where $a$ is an arbitrary apex in $A_{X}$. If $b_{1}$ and $b_{2}$ belong to a common path $P\in\mathcal{P}$ then, on the induced path between them, let $b^{\prime}_{1}$ be the neighbour of $b_{1}$ and $b^{\prime}_{2}$ that of $b_{2}$. If not, let $b^{\prime}_{1}=b_{1}+2$ and $b^{\prime}_{2}=b_{2}+2$.
@@
Figure 1. Salvaging a “failed” apex $a$. Upon offer of suitable odd points $b_{1}$ and $b_{2}$, hyperedge $(b_{1},a,b_{2})$ is introduced (dashed line) and added to $\mathcal{C}$. The good base pairs $\left\{b_{1},b_{1}^{\prime}\right\}$ and $\left\{b_{2},b_{2}^{\prime}\right\}$ (dotted lines) are deleted from $D$. (As a failed apex, $a$ was already absent from $D$.) This converts two paths in $\mathcal{P}$ (the figure’s upper and lower lines) into three, or one path into two if $b_{1}$ and $b_{2}$ lie on a common path (if in the figure, $t_{1}=t_{2}$).
@@
The action incorporates the failed apex $a$ into a hyperedge, at the expense of removing two good base pairs $(b_{1},b^{\prime}_{1})$ and $(b_{2},b^{\prime}_{2})$ from $D$, and turning two paths in $\mathcal{P}$ into three (as shown in the figure, when $b_{1}$ and $b_{2}$ are in distinct paths) or one into two (when $b_{1}$ and $b_{2}$ are in a common path). The action decreases the outdegree of any node in $D$ by at most 1: it is only $(b_{1},b^{\prime}_{1})$ and $(b_{2},b^{\prime}_{2})$ that are deleted, by condition (H2a-C2) neither was in $Q$ therefore neither has a double inedge, and by condition (H2a-C3) they have no common inneighbour.
In any trial, the conditions are likely to be satisfied. For (H2a-C1), there is probability $O(1/n)$ that $b_{1}$ and $b_{2}$ are path neighbours. For (H2a-C2), $\mathcal{P}$ has $(1/2-O(\epsilon_{D}))n$ edges on its $n/2$ points, so most points have two edges in $\mathcal{P}$ (so it is likely $b_{1}$ and $b_{2}$ do) while only $O(\epsilon_{Q})n$ points are in $Q$ (so it is unlikely that $b_{1}$ or $b_{2}$ are). Finally, (H2a-C3) follows from 8. Since all of the handful of failure events have probability $O(\epsilon_{D}+\epsilon_{Q})$, so does their union.
@@ 5.2.2. Phase 2b
Figure 2. Patching together base-pair paths. Suppose a previous offer included an apex $a$ and base point $b$, to which we added some path end $s_{1}$ to propose hyperedge $(b,a,s_{1})$. If by coincidence the current offer includes base point $b+2$ and an apex $a^{\prime}$, then, for some the opposite path end $t_{1}$, propose hyperedge $(b+2,a^{\prime},t_{1})$. Accepting this pair of proposals means adding the (dashed) path edges $\left\{b,s_{1}\right\}$ and $\left\{b+2,t_{1}\right\}$ and deleting the (dotted) edge $\left\{b,b+2\right\}$, turning two $\mathcal{P}$ paths into one. Correspondingly, we accept the two hyperedges into $\mathcal{C}$, and delete the base pair $\left\{b,b+2\right\}$ from $B$.
@@
We partition the $n/2$ odd points into $n/4$ “cell” pairs, $\left\{1,3\right\},\allowbreak\left\{5,7\right\},\allowbreak\ldots,\allowbreak\left\{n-3,n-1\right\}$; these pairs are fixed for all rounds. Hyperedges added in a round are used only within that round; except for those used in actions, they are ignored in future rounds and can be thought of as deleted.
@@
Consider only offers of the form $(b,a)$, odd and even, discarding others. Also discard any offer where $b$ was offered earlier in this round. Assume then that this is the first time $b$ is offered. If the cell partner $b^{\prime}$ of $b$ also has not been offered, complete this offer to $(b,a,t)$ where $t$ is a random terminal, and add $(b,a,t)$ to $\mathcal{H}$. Consider the cell “seeded”. If the cell partner $b^{\prime}$ was previously seen (the cell was seeded), then some hyperedge $(b^{\prime},a^{\prime},t_{i})$ is in $\mathcal{H}$, in which case complete the current offer to $(b,a,s_{i})$. Consider the cell of $b$ “filled”, and “attempt action”. Attempting action means taking action (see below) if the following conditions are satisfied:
@@
Although the offers arrive successively (and the round is terminated once $(1/10)\ell$ actions are taken), consider the round’s full $10f$ offers. Each offer has probability just over $1/2$ of being of the requisite form $(b,a)$. Each one of that form has equal probability of $b$ being a “left” or “right” element of a cell, i.e., of $b\equiv 1\pmod{4}$ or $b\equiv 3\pmod{4}$. In $10f$ offers, with probability $1-\exp(-\Omega(n))$, there will be at least $2.4f$ each lefts and rights; we assume henceforth that this is so.
@@
When a new cell is filled, the conditions for action are satisfied with probability at least $2/10$, assuming that the number of paths in $\mathcal{P}$ is at most $(9/10)\ell$.
@@
The algorithm terminates if the number of attempted actions — cells filled — exceeds $f^{2}/n$. Thus, the probability that a random cell $X$ was previously filled is $\Pr(\overline{F})\leq(f^{2}/n)/(n/4)=4(f/n)^{2}\leq 2/100^{2}$ by (5.5). So, $\Pr(F)\geq 0.99$. It follows that $\Pr(E\mid F)\geq\Pr(E)/\Pr(F)\geq 2/10$. ∎
@@
$\frac{\frac{1}{2}f^{2}/n}{\ell}$ $=\frac{n^{3/2}\ell^{1/2}}{2n\ell}=\frac{1}{2}(n/\ell)^{1/2}\geq\frac{1}{2}(1/\epsilon_{D})^{1/2}>1.$
@@ 5.3. Phase 3
To avoid excessive notation, consider $s=4$ for example. Take base pairs $\left\{1,4\right\},\left\{4,7\right\},\left\{7,10\right\},\ldots$, and apex nodes (no longer single points) $\left\{2,3\right\},\left\{5,6\right\},\left\{8,9\right\},\ldots$. The generalisation to other $s$ is clear. We will see that it is easy to keep each apex node set intact, so that it behaves just like an apex point in the $s=3$ case.
@@ Acknowledgements
Greg is grateful to Alan Frieze for discussions of Hamilton cycles, and to Olaf Parczyk for extensive discussions of the semirandom hypergraph process.

### v2

@@ 5. Loose Hamilton cycles
The next steps are structurally the same as for matchings in Section 4. As in Section 4, in the 1-offer model we could (exactly as in Section 4) make a 2-out graph on $A_{0}$ and $B_{0}$; it would a.a.s. contain a perfect matching; and that auxiliary graph matching corresponds to a loose Hamilton cycle in the hypergraph. In the 2-offer model it is not so easy.
We will construct a directed bipartite graph on $A_{0}$ and $B_{0}$, but it will not have a perfect matching because a small number of nodes fail to have sufficiently high degree (and cause other nodes to “fail” as well). We deal with the failed nodes first, and in doing so add certain hyperedges, irrevocably, to our cycle-in-the-making $\mathcal{C}$. ($\mathcal{C}$ plays the role that $\mathcal{M}$ did for hypergraph matchings.) When the failed nodes are dealt with, most of our bipartite graph will remain, and what remains will have a perfect matching. The matching edges correspond to hyperedges in $\mathcal{H}$, and they complete $\mathcal{C}$ to a loose Hamilton cycle.
@@ 5.2. Phase 2
We introduce an additional structure, $\mathcal{P}$, central to the analysis. $\mathcal{P}$ is a graph on the odd points, consisting of vertex-disjoint paths. In particular, a path in $\mathcal{P}$ can be an isolated vertex. Initially, the edges in $\mathcal{P}$ are precisely the base pairs $B_{1}$. Eventually we will make $\mathcal{P}$ a single path and then a Hamilton cycle. Always, the edges in $\mathcal{P}$ are of two types: every hyperedge $e=(b_{1},a,b_{2})\in\mathcal{H}$ contributes a $\mathcal{P}$-edge $(b_{1},b_{2})$ (and the apexes $a$ of these will all be distinct); and every base pair in $B$ is an edge in $\mathcal{P}$ (and later it will be possible to perfectly match all these base pairs and the remaining apexes, in $D$). When $\mathcal{P}$ is a Hamilton cycle, this ensures that there is a corresponding loose hypergraph cycle $\mathcal{C}$ as desired.
In this phase we will again use a set $Q$ of nodes of $D$ “blocked” from use. Initially we will set $Q=Q_{0}$ to consist of any nodes whose inedges include any double edge; recall from 7 that $\left|{Q_{0}}\right|=o(n)$. Initially, $\mathcal{P}$ has $b_{x}$ components. Let us describe the phase just enough to count the number of nodes it will delete from each part of $D$, both to bound $\left|{Q}\right|$ and to confirm that, going in to Phase 3, $\left|{A}\right|=\left|{B}\right|$.
In Phase 2a, each action (see Figure 1) will assign an apex in $A_{X}$ to a hyperedge (along with two base points from different base pairs) and place it in $\mathcal{C}$, increasing the number of paths in $\mathcal{P}$ by 1, and deleting 2 base pairs from $B$. Phase 2a takes $a_{x}$ actions to resolve all the failed apexes, resulting in a total of $2a_{x}$ deletions from the set of base pairs in $B$, so that $\left|{B_{0}\setminus B}\right|=b_{x}+2a_{x}$, and there are $b_{x}+a_{x}$ components in $\mathcal{P}$. The part $A$ is unchanged, with $\left|{A_{0}\setminus A}\right|=a_{x}$.
@@
We can make the blocked set $Q$ arbitrarily small by choosing $\epsilon_{D}$ sufficiently small, as we now show. For any given $\epsilon_{Q}>0$, let Lemma 9’s $\eta$ equal $\epsilon_{Q}/k$, and take the lemma’s corresponding $\delta$ to determine $4\epsilon_{D}$ (choosing $\epsilon_{D}$ smaller if required elsewhere). Let $S$ be the set of deleted nodes. Then, Lemma 9 ensures that the deleted set’s inneighbourhood has size $\left|{N^{-}(S)}\right|\leq\eta\>N\leq(\epsilon_{Q}/k)n$. Since $D$ is $k$-out, the inneighbourhood’s outneighbourhood — which is to say, the rest of $Q$ beyond $Q_{0}$ — has size
$\left|{Q\setminus Q_{0}}\right|=\left|{N^{+}(N^{-}(S))}\right|\leq k(\epsilon_{Q}/k)n=\epsilon_{Q}n.$
@@ 5.2.1. Phase 2a
Upon offer of odd points $b_{1}$ and $b_{2}$, we propose a hyperedge $e=(b_{1},a,b_{2})$, where $a$ is an arbitrary apex in $A_{X}$. If $b_{1}$ and $b_{2}$ belong to different paths in $\mathcal{P}$ then let $b^{\prime}_{1}=b_{1}+2$ and $b^{\prime}_{2}=b_{2}+2$. If they belong to a common path $P$ in $\mathcal{P}$ then, on the induced path between them, let $b^{\prime}_{1}$ be the neighbour of $b_{1}$ and $b^{\prime}_{2}$ that of $b_{2}$. (In the common-path case, typically $b_{1}^{\prime}=b_{1}\pm 2$ and $(b_{1},b_{1}^{\prime})\in B$, and likewise for $(b_{2},b_{2}^{\prime})$; more on this soon.)
@@
Figure 1. Salvaging a “failed” apex $a$ in Phase 2a. Upon offer of suitable odd points $b_{1}$ and $b_{2}$, hyperedge $(b_{1},a,b_{2})$ is introduced (dashed line) and added to $\mathcal{C}$. The good base pairs $\left\{b_{1},b_{1}^{\prime}\right\}$ and $\left\{b_{2},b_{2}^{\prime}\right\}$ (dotted lines) are deleted from $D$. (As a failed apex, $a$ was already absent from $D$.) This converts two paths in $\mathcal{P}$ (the figure’s upper and lower lines) into three, or one path into two if $b_{1}$ and $b_{2}$ lie on a common path (if in the figure, $t_{1}=t_{2}$).
@@
The action incorporates the failed apex $a$ into a hyperedge, at the expense of removing two good base pairs $(b_{1},b^{\prime}_{1})$ and $(b_{2},b^{\prime}_{2})$ from $D$, and turning two paths in $\mathcal{P}$ into three (as shown in the figure, when $b_{1}$ and $b_{2}$ are in distinct paths) or one into two (when $b_{1}$ and $b_{2}$ are in a common path). Every pair in $B$ should be an edge of $\mathcal{P}$, and since the action deletes $(b_{1},b^{\prime}_{1})$ and $(b_{2},b^{\prime}_{2})$ from $B$ as well as from $\mathcal{P}$, this remains the case. The action decreases the outdegree of any node in $D$ by at most 1: it is only $(b_{1},b^{\prime}_{1})$ and $(b_{2},b^{\prime}_{2})$ that are deleted, by condition (H2a-C2) neither was in $Q$ therefore neither has a double inedge, and by condition (H2a-C3) they have no common inneighbour.
In any trial, the conditions are likely to be satisfied. For (H2a-C1), there is probability $O(1/n)$ that $b_{1}$ and $b_{2}$ are path neighbours. For (H2a-C2), $b_{1}^{\prime}$ is either $b_{1}+2$ or a $\mathcal{P}$-neighbour of $b_{1}$, but it is likely that the $\mathcal{P}$-neighbours of $b_{1}$ are just $b_{1}\pm 2$, because the $(1/2-O(\epsilon_{D}))n$ edges in $\mathcal{P}$ include the $(1/2-O(\epsilon_{D}))n$ pairs in $B$. And if $b_{1}^{\prime}=b_{1}\pm 2$, it is likely that $(b_{1},b_{1}^{\prime})\in B\setminus Q$, since $B$ includes $(1/2-O(\epsilon_{D}))n$ pairs of this form while $\left|{Q}\right|=O(\epsilon_{Q})n$. The same holds for $b_{2}^{\prime}$. Finally, (H2a-C3) follows from 8. Since all of the handful of failure events have probability $O(\epsilon_{D}+\epsilon_{Q})$, so does their union.
@@ 5.2.2. Phase 2b
Figure 2. Patching together base-pair paths in Phase 2b. Suppose a previous offer included an apex $a$ and base point $b$, to which we added some path end $s_{1}$ to propose hyperedge $(b,a,s_{1})$. If by coincidence the current offer includes base point $b+2$ and an apex $a^{\prime}$, then, for some the opposite path end $t_{1}$, propose hyperedge $(b+2,a^{\prime},t_{1})$. Accepting this pair of proposals means adding the (dashed) path edges $\left\{b,s_{1}\right\}$ and $\left\{b+2,t_{1}\right\}$ and deleting the (dotted) edge $\left\{b,b+2\right\}$, turning two $\mathcal{P}$ paths into one. Correspondingly, we accept the two hyperedges into $\mathcal{C}$, and delete the base pair $\left\{b,b+2\right\}$ from $B$.
@@
We partition the $n/2$ odd points into $\left\lfloor n/4\right\rfloor$ disjoint “cell” pairs $\left\{1,3\right\},\allowbreak\left\{5,7\right\},\left\{9,11\right\},\allowbreak\ldots$. (If $n\equiv 2\pmod{4}$ there is also an odd-man-out singleton $\left\{n-1\right\}$; it will never be used, is merely a pesky detail, and will only be mentioned once again.) These cells are fixed for all rounds. Hyperedges added in a round are used only within that round; except for those used in actions, they are ignored in future rounds and can be thought of as deleted.
@@
Consider only offers of the form $(b,a)$, $b$ odd (and not the pesky odd-one-out singleton, if any) and $a$ even, discarding others. Also discard any offer where $b$ was offered earlier in this round. Assume then that this is the first time $b$ is offered. If the cell partner $b^{\prime}$ of $b$ also has not been offered, complete this offer to $(b,a,t)$ where $t$ is a random terminal, and add $(b,a,t)$ to $\mathcal{H}$. Consider the cell “seeded”. If the cell partner $b^{\prime}$ was previously seen (the cell was seeded), then some hyperedge $(b^{\prime},a^{\prime},t_{i})$ is in $\mathcal{H}$, in which case complete the current offer to $(b,a,s_{i})$. Consider the cell of $b$ “filled”, and “attempt action”. Attempting action means taking action (see below) if the following conditions are satisfied:
@@
Although the offers arrive successively (and the round is terminated once $(1/10)\ell$ actions are taken), consider the round’s full $10f$ offers. Each offer has probability asymptotically $1/2$ of being of the requisite form $(b,a)$. Each one of that form has equal probability of $b$ being a “left” or “right” element of a cell, i.e., of $b\equiv 1\pmod{4}$ or $b\equiv 3\pmod{4}$. In $10f$ offers, with probability $1-\exp(-\Omega(n))$, there will be at least $2.4f$ each lefts and rights; we assume henceforth that this is so.
@@
When a new cell is filled, the conditions for action are satisfied with probability at least $2/10$, assuming that the number of paths in $\mathcal{P}$ is at least $(9/10)\ell$.
@@
The algorithm terminates if the number of attempted actions — cells filled — exceeds $f^{2}/n$. Thus, the probability that a random cell $X$ was previously filled is $\Pr(\overline{F})\leq(f^{2}/n)/(n/4)=4(f/n)^{2}\leq 4/100^{2}$ by (5.5). So, $\Pr(F)\geq 0.99$. It follows that $\Pr(E\mid F)\geq\Pr(E\cap F)\geq\Pr(E)-\Pr(\bar{F})\geq 2/10$. ∎
@@
$\frac{\frac{1}{10}f^{2}/n}{\ell}$ $=\frac{n^{3/2}\ell^{1/2}}{10n\ell}=\frac{1}{10}\,(n/\ell)^{1/2}\geq\frac{1}{10}(1/\epsilon_{D})^{1/2}>1,$
for $\epsilon_{D}$ sufficiently small.
@@ 5.3. Phase 3
To avoid excessive notation, consider $s=4$ for example. Take base pairs $\left\{1,4\right\},\allowbreak\left\{4,7\right\},\allowbreak\left\{7,10\right\},\ldots$, and apex nodes (no longer single points) $\left\{2,3\right\},\left\{5,6\right\},\left\{8,9\right\},\ldots$. The generalisation to other $s$ is clear. We will see that it is easy to keep each apex node set intact, so that it behaves just like an apex point in the $s=3$ case.
@@ Acknowledgements
Greg is grateful to Alan Frieze for guidance on graph Hamilton cycles, to Olaf Parczyk for extensive discussions on semirandom hypergraphs, and to Warach Veeranonchai for a careful reading resulting in several improvements to the manuscript.